# A Collaborative Model-driven Network for MRI Reconstruction

### v1

@@ Abstract
Magnetic resonance imaging (MRI) is a vital medical imaging modality, but its development has been limited by prolonged scanning time. Deep learning (DL)-based methods, which build neural networks to reconstruct MR images from undersampled raw data, can reliably address this problem. Among these methods, model-driven DL methods incorporate different prior knowledge into deep networks, thereby narrowing the solution space and achieving better results. However, the complementarity among different prior knowledge has not been thoroughly explored. Most of the existing model-driven networks simply stack unrolled cascades to mimic iterative solution steps, which are inefficient and their performances are suboptimal. To optimize the conventional network structure, we propose a collaborative model-driven network. In the network, each unrolled cascade comprised three parts: model-driven subnetworks, attention modules, and correction modules. The attention modules can learn to enhance the areas of expertise for each subnetwork, and the correction modules can compensate for new errors caused by the attention modules. The optimized intermediate results are fed into the next cascade for better convergence. Experimental results on multiple sequences showed significant improvements in the final results without additional computational complexity. Moreover, the proposed model-driven network design strategy can be easily applied to other model-driven methods to improve their performances.
{IEEEkeywords}
attention mechanism, low-rankness, MRI reconstruction, model-driven, sparsity.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is an irreplaceable imaging modality [1], which has been widely used in clinical scenarios. However, the acquisition of full-sampled raw k-space data is inherently time-consuming. This aggravates patient discomfort and produces artifacts on MR images, limiting the development and application of MR imaging.
To address this problem, parallel imaging (PI)[2], which uses multiple receiving coils to acquire MR signals, has been widely employed in clinical scenarios. PI methods are classified into two main categories: elimination of aliasing artifacts in the image domain, which is represented by the SENSE[3] algorithm; and interpolation in the k-space domain, which is represented by the GRAPPA[4] algorithm. In addition to PI, compressed sensing (CS) theory [5, 6], which is highly compatible with the MR imaging background, has also been applied to MR image reconstruction [7, 8]. The CS theory enables the reconstruction of MR images from highly undersampled k-space data, which violates the Nyquist-Shannon sampling theorem. The sparsity of the k-space data in the transform domains is typically exploited to narrow the solution space. Iterative optimization algorithms are then leveraged to approach the target MR images.
Figure 1: (a) conventional model-driven network structure. (b) common modification in model-driven networks. (c) our proposed network structure.
@@
Conventional approaches are limited by issues such as long computation times, low acceleration rates, and suboptimal performance [9, 10]. Considering the significant progress made by deep learning (DL) methods in computer vision tasks, DL-based methods have been successfully introduced into the MRI reconstruction field, which can be broadly categorized into data-driven and model-driven methods [11]. Data-driven methods [12, 13, 14, 15, 16, 17] use deep networks to build a pipeline between the inputs (undersampled k-space data or aliased images) and outputs (fully sampled data or artifact free images). In comparison, model-driven methods [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] can be viewed as combinations of conventional and data-driven methods. Prior knowledge is integrated into model-driven networks, whereas deep-learning methods can learn more complex feature representations through learning processes. More specifically, the solution steps of conventional algorithms are unrolled into deep networks, as shown in Fig. 1 (a). For example, Zhang [19] et al. proposed ISTA-Net that unfolds the update steps of ISTA [30] into a deep network. The sparsity transform and its inverse version in ISTA were learned in the network, and a loss function was designed to fulfil the symmetry constraint. In contrast, Duan [20] et al. developed a variable-splitting method to solve the objective function and incorporated the corresponding three-step update process into the network. Sriram [21] et al. extended the Variational Network [22] and fulfiled the update steps in k-space by learning end-to-end. Yiasemis [23] et al. explored the applied Convolutional Recurrent Neural Networks (CRNNs) to an unrolled network to solve the inverse problem.
@@
A common way to optimize model-driven methods is to incorporate different regularizers into the objective functions to narrow the solution space more precisely (Fig. 1 (b)). Some methods [46, 47, 48, 49] simultaneously incorporated prior knowledge, such as sparsity and low rankness, in dynamic MRI considering the image characteristics. In DL-based reconstruction methods [35, 41], objective functions designed with sparsity- and low-rankness-based regularizers were used to built a network based on a corresponding iterative algorithm. In addition to explicit prior knowledge, some methods also proposed to impose constraints from learned sensitivity maps for accurate reconstructions [27, 50]. However, the existing methods have not explored the way of fully exploiting the complementarity of different types of priors. Most model-driven methods are designed to stack unrolled cascades, and different priors can then aid in network training in an alternative form. We believe the network design is inefficient, and their reconstruction performance is suboptimal. Instead, we aim to propose a parallel network structure that different branches in the network can simultaneously contribute to the training process to achieve better results. Therefore, we designed a new strategy for designing model-driven networks (Fig. 1 (c)). Our objective is to find the areas of expertise of the subnetworks based on different priors to further optimize intermediate results and seek better convergence. Alternatively, from a macro-perspective, we can easily determine the global optimum by appropriately integrating different priors into the network.
The proposed network is called a collaborative model-driven network (CMD-Net). In the proposed network, subnetworks based on sparsity and low rankness are effectively integrated using an attention module to achieve better performances. Correction modules are connected to the optimized weighted images, providing considerable flexibility to compensate for newly introduced errors. Data consistency modules were used to maintain consistency with the known sampling information. Although the subnetworks are established by stacking the unrolled cascades of iterative algorithms, the optimized intermediate results from the last cascade are used to replace the inputs of the next cascade to facilitate network convergence. Moreover, the intermediate results from different modules are directly input into the loss function to constrain their functions. The experimental results show that the proposed method outperforms one-prior and traditional multiprior-based networks. Compared with the state-of-the-art methods, our method also yields better results. Furthermore, the proposed method is compatible with other model-driven networks having different priors or objective functions. The experimental results show that using the proposed strategy, the performance of different model-driven networks can be improved without increasing computational complexity.
@@
    We propose a collaborative model-driven MRI reconstruction network design strategy, which effectively integrate sub-networks based on different priors.
@@
    Experimental results on multiple sequences show that the proposed method can achieve better performance without extra computational complexity.
@@
    The proposed network design strategy can be easily applied to other model-driven network and significantly improve their performance.
@@ 2 Problem Formulation
in sparsity-based algorithms. $\Phi$ is an appropriate transformation field for ensuring the sparsity of the raw data. In this study, we aimed to demonstrate the generality and efficiency of the proposed network design strategy. Therefore, we adopted an ISTA-based backbone to solve (7), which iterates between the following two steps:
@@
For a low-rankess-based method, (6) can be expressed as
@@
$\hat{x}^{k}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x^{k-1}-y\|_{2}^{2}+\lambda_{2}\|\tau(x^{k-1})Q^{k-1}\|_{F}^{2},$ (13)
@@
$\hat{x}^{k}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x^{k-1}-y\|_{2}^{2}+\lambda_{2}\|\mathcal{D}(Q^{k-1})x^{k-1}\|_{F}^{2},$ (16)
and inspired by previous studies [52, 53, 41], we solve (16) by iterating the following two steps:
@@ 3 Method
Figure 2: (a): the structure of the proposed network. (b): the structure of the sparsity based sub-network ($N_{s}$ in (a)). (c): the structure of the low-rank based sub-network ($N_{l}$ in (a)). (d): the structure of the attention modules (the yellow blocks in (a)), each of which include a channel combination block (CCB), an attention block (AB), and a weighting block (WB). (e): the structure of the correction modules (the blue blocks in (a)), each of which includes a regularization block (RB) and a data consistency block.
@@ 3.1 Network Structure
The proposed network structure and its components are illustrated in Fig. 2. Fig. 2 (a) shows the overall network structure. In the network training process, the undersampled k-space data $y\in\mathbb{C}^{NC}$ as well as the corresponding sensitivity maps $\mathcal{S}\in\mathbb{C}^{NC}$ and masks $\mathcal{P}\in\mathbb{R}^{N}$ are fed to the network. MR images are stored using complex values, and some studies have proposed complex convolutions [13] to achieve better results; however, this is not our focus. Therefore, two channels were used to store the real and imaginary parts in this study, and $y\in\mathbb{C}^{NC}$, $\mathcal{S}\in\mathbb{C}^{NC}$ were transferred to $y\in\mathbb{R}^{2NC}$, $\mathcal{S}\in\mathbb{R}^{2NC}$ before they are fed into the network. Besides, the coil-combined image $x_{0}$ is obtained from the measurement $y$ and fed to the network. Then, several iteration blocks were cascaded to form the main body, and each block is referred to as a cascade. Through stacked cascades, the network finally outputs the reconstructed MR images. The final and intermediate results were fed into the customized loss function introduced in Section 3.5. Each cascade comprised three parts: model-driven subnetworks (green blocks), attention-based fusion modules (yellow blocks), and correction modules (blue blocks). Fig. 2 (b)–(e) present the detailed structures of the modules, which are introduced in the following subsections.
3.2 Model-driven Sub-networks
@@
$N_{l}$ in each cascade is constructed by unrolling (17) and (18) into the network. For (17), essentially $(\mathcal{D}(Q))^{H}\mathcal{D}(Q)x$ can be regarded as a projection acting on $x$, which is determined by $Q$, and $Q$ at each iteration will also be updated according to the status of $x$. Therefore, we relaxed the constraint on (17) and used U-Net to learn the mapping directly, as shown in Fig. 2 (c). By doing so, we avoided a complicated calculation process, and more accurate intermediate variables can be learned directly from the data. We used $N_{lu}$ to represent U-Net, and thus, we have
@@
Notably, the proposed model-driven subnetworks differ from another previously proposed method [41] that also involves sparsity and low-rank priors. First, they proposed decoupling the 2D Fourier transform into two 1D processes, and then obtained much more data to train the network, while greatly reducing the computational complexity. However, our purpose was to design a universal model-driven network that can be easily generalized. Therefore, we still adopted a 2D reconstruction process. More importantly, we do not place both regularizers into a single objective function, which is
The corresponding model-driven network was built based on the solution steps. As an alternative, we place different regularizers based on different priors into two objective functions (i.e., (7) and (10)). Hence, two subnetworks are built. In our opinion, the iterative steps for solving (22) essentially incorporate the information learned from regularizers into the training step in an alternating form. For comparison, we aimed to effectively integrate the intermediate results from different subnetworks and make different branches simultaneously contribute to the training process. By doing so, different subnetworks can help each other achieve faster convergence. We conducted experiments in Section 4.3 to compare the two aforementioned methods. Then, we propose an attention module to organically integrate the subnetworks, which is introduced in the next subsection.
Refer to caption coronal PD 4x
Figure 3: From left to right: a fully sampled MR image, a zero-filled undersampled MR image and the corresponding sampling mask. The first row: 4-fold acceleration. The second row: 6-fold acceleration.
3.3 Attention Module
The attention module (ATM) includes all the yellow blocks in Fig. 2; the details are shown in Fig. 2 (d). The two inputs are $x_{s}^{k}\in\mathbb{R}^{2N}$ and $x_{l}^{k}\in\mathbb{R}^{2N}$, which are the outputs of $N_{s}$ and $N_{l}$ in the $k$th cascade. An ATM aims to find better intermediate results for $x_{s}^{k}$ and $x_{l}^{k}$ for each data point. To achieve this, we first propose a channel combination block (CCB) to concatenate $x_{s}^{k}$ and $x_{l}^{k}$ on the channel dimensions, and we denote the combined result as $x_{c}^{k}\in\mathbb{R}^{4N}$. Then, $x_{c}^{k}$ is fed into an element-wise attention block comprising two convolutional layers, each of which connects to an activation layer. We chose the ReLU and sigmoid functions. After the four-layer attention block, we obtained an attention map $M^{k}\in\mathbb{R}^{N}$, where each element in $M^{k}$ lies between zero and one. We transfered $x_{s}^{k}$ and $x_{l}^{k}$ back to complex values (i.e., $x_{s}^{k},x_{l}^{k}\in\mathbb{R}^{2N}\rightarrow x_{s}^{k},x_{l}^{k}\in\mathbb{C}^{N}$) and then created a Hadamard product between $M^{k}$ and $x_{s}^{k}$. $x_{l}^{k}$ is similarly multiplied by $\textbf{J}-M^{k}$. $\textbf{J}\in\mathbb{R}^{N}$ is a matrix filled with one. We used $N_{a}$ to represent the weighting block, and the weighted output is given by
$x^{k}_{a}=N_{a}(x_{s}^{k},x_{l}^{k})=x^{k}_{s}\odot M^{k}+x^{k}_{l}\odot(\textbf{J}-M^{k})$ (23)
Therefore, the closer the number in $M^{k}$ is to 1, the higher the value in $x^{k}_{s}$ at the same place occupied in $x^{k}_{a}$. To directly supervise the attention module, $x_{a}^{k}$ explicitly contributes to the loss function. Therefore, the attention module assigns more weights to the subnetwork on the data points that perform better.
3.4 Correction Module
Through ATMs, the data points on MR images reconstructed by proper subnetworks can have larger weights, and the gap between intermediate outputs and targets can be narrowed. However, this process may introduce new errors owing to inaccurate weights. Therefore, we propose a correction module to revise the errors caused by ATMs. Each correction module consists of a regularizer block (RB) and a data consistency (DC) block. RBs are designed to correct the intermediate results in the image domain and DC blocks in k-space, so that we can fully exploit the complementarity between cross-domain information.
First, a RB block is designed to enhance the network learning ability and provide more flexibility, which is

### v2

@@ Abstract
Deep learning (DL)-based methods offer a promising solution to reduce the prolonged scanning time in magnetic resonance imaging (MRI). While model-driven DL methods have demonstrated convincing results by incorporating prior knowledge into deep networks, further exploration is needed to optimize the integration of diverse priors.. Existing model-driven networks typically utilize linearly stacked unrolled cascades to mimic iterative solution steps in optimization algorithms. However, this approach needs to find a balance between different prior-based regularizers during training, resulting in slower convergence and suboptimal reconstructions. To overcome the limitations, we propose a collaborative model-driven network to maximally exploit the complementarity of different regularizers. We design attention modules to learn both the relative confidence (RC) and overall confidence (OC) for the intermediate reconstructions (IRs) generated by different prior-based subnetworks. RC assigns more weight to the areas of expertise of the subnetworks, enabling precise element-wise collaboration. We design correction modules to tackle bottleneck scenarios where both subnetworks exhibit low accuracy, and they further optimize the IRs based on OC maps. IRs across various stages are concatenated and fed to the attention modules to build robust and accurate confidence maps. Experimental results on multiple datasets showed significant improvements in the final results without additional computational costs. Moreover, the proposed model-driven network design strategy can be conveniently applied to various model-driven methods to improve their performance.
keywords:
MRI reconstruction , model-driven , attention mechanism , sparsity , low-rankness
journal: Biomedical Signal Processing and Control
[…]
@@ 1 Introduction
Magnetic resonance imaging (MRI) is an irreplaceable imaging modality widely employed in clinical scenarios. However, the acquisition of full-sampled raw k-space data is inherently time-consuming, which aggravates patient discomfort and produces artifacts on MR images, limiting the development and application of MR imaging [1].
To address the problem, sub-Nyquist sampling rates are routinely adopted to reduce scanning time, and techniques including parallel imaging (PI) [2] and compressed sensing (CS) [3, 4] enable the reconstruction of MR images from undersampled data. The PI methods [5, 6] exploit the redundancy among k-space data, but the image quality deteriorates at high acceleration factors. In CS-MRI [7, 8], prior knowledge such as sparsity of the k-space data in transform domains is typically exploited to narrow the solution space. Iterative optimization algorithms are then designed to approach the target MR images. However, CS-based methods still face limitations such as reliance on hand-crafted tuning and prolonged iteration processes. [9, 10].
With the notable advancements achieved by deep learning (DL) methods in computer vision tasks, DL-based methods have been successfully introduced into the MRI reconstruction field, which can be broadly categorized into data-driven and model-driven methods [11]. The end-to-end data-driven methods [12, 13, 14, 15, 16, 17, 18] use deep networks to build a pipeline between the inputs (undersampled k-space data or aliased images) and outputs (fully sampled data or artifact-free images), which require a large amount of training data. Another broad family of data-driven methods focuses on distribution learning. Diffusion models including denoising diffusion probablistic models (DDPM) and score matching with Langevin dynamics (SMLD) have demonstrated convincing results in accelerated MRI. However, the reverse process in diffusion models is time-consuming and not reliable enough for medical imaging purposes.
@@
In comparison, model-driven methods [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] can be viewed as combinations of conventional and data-driven methods. Prior knowledge is integrated into model-driven networks, whereas deep networks can learn more complex feature representations through learning processes. More specifically, the solution steps of conventional algorithms are unrolled into deep networks. For example, Zhang [20] et al. proposed ISTA-Net that unfolds the update steps of ISTA [32] into a deep network. The sparsity transform and its inverse version in ISTA were learned in the network, and a loss function was designed to fulfil the symmetry constraint. Duan [21] et al. developed a variable-splitting method to solve the objective function and incorporated the corresponding three-step update process into the network. Sriram [22] et al. extended the Variational Network [23] and fulfilled the update steps in k-space by learning end-to-end. Yiasemis [24] et al. employed Convolutional Recurrent Neural Networks (CRNNs) within an unrolled network to address the inverse problem. In [33], the optimization process was established using half-quadratic variable splitting method, and a dilated-convolution model was employed to the predict Lagrange multipliers.
@@
A common way to optimize model-driven methods is to incorporate different prior-based regularizers into the objective functions to narrow the solution space more precisely and better avoid falling into local minima. For example, some methods [49, 50, 51, 52] simultaneously incorporated prior knowledge, such as sparsity and low rankness, in dynamic MRI considering the image characteristics. In DL-based reconstruction methods [38, 44], objective functions incorporating sparsity- and low-rankness-based regularizers were used to build a network based on a corresponding iterative algorithm. In addition to explicit prior knowledge, some methods also proposed to impose constraints from learned sensitivity maps for accurate reconstructions [28, 53]. To unroll the optimization steps of a multi-prior-based problem into model-driven networks, unrolled blocks are linearly stacked to alternately introduce constraints from different priors. However, the network needs to find a balance between different prior-based regularizers, leading to slower convergence and consequently suboptimal outcomes. From a more microscopic perspective, different regularization-based blocks specialize in different positions of the target images, resulting in varying speeds to approach the target values for each pixel. Ideally, selecting the most accurate reconstructed results in an element-wise manner in each cascade can effectively accelerate network convergence. This, in turn, allows for more precise intermediate results to be fed into subsequent cascades, increasing the likelihood of approaching the target images with more accurate inputs for the following cascades.
To this end, we propose a versatile collaborative network that different branches based on different priors can simultaneously contribute to the training process to achieve better results. The proposed network is called a collaborative model-driven network (CMD-Net). Our objective is to find the areas of expertise of the subnetworks based on different priors to produce more accurate results for each pixel and seek better convergence. We design three modules in each unrolled cascade within the network: model-driven subnetworks (MNs), attention modules (AMs), and correction modules (CMs). In the proposed network, MNs based on sparsity and low-rankness produce intermediate reconstructions (IRs) in each cascade. With different prior knowledge, the IRs show varying performance in different positions. To fully exploit them, we design an AM to learn the relative confidence (RC) and overall confidence (OC) for these IRs. IRs across various stages are concatenated and fed to the AMs to build robust and precise mappings to the confidence maps. The IRs are fused with the RC maps, which assign more weights to the areas of expertise of the corresponding subnetworks in an element-wise manner. Considering that when both subnetworks exhibit low accuracy in specific locations, the weighting process cannot effectively improve the results, we further propose a CM. Unlike the subnetworks, the mapping in CMs is directly learned from data without incorporating explicit priors, providing considerable flexibility to compensate for errors brought by the IRs. When the IRs have sufficient high OC values, which means there is high confidence in accepting the RC-fused IRs as the final ones, the correction modules will operate within a relatively small correction range. Otherwise, more weights are assigned to the correction results to further optimize the fused IRs. Data consistency (DC) blocks in CMs were used to maintain consistency with the known sampling data. The optimized IRs from each cascade are used to replace the inputs of the next cascade to accelerate network convergence. Moreover, the results from each cascade are directly fed into the loss function to constrain their functions, and inaccurate confidence maps will result in higher losses. We adopt floating-point operations (FLOPs) to compare the computational costs of different networks. The experimental results show that the proposed method clearly outperforms the networks that linearly stack unrolled cascades. Compared with the state-of-the-art methods on multiple datasets, our method also yields better results. Furthermore, the proposed method is compatible to optimize the existing model-driven networks based on single- or multi-prior. The experimental results show that using the proposed strategy, the performance of different model-driven networks can be improved without increasing computational costs.
@@
    We propose a novel collaborative model-driven network design strategy for MRI reconstruction, and the proposed attention-based fusion modules and correction modules in each cascade can effectively maximize the contributions of different prior knowledge based subnetworks to achieve fast and robust convergence.
@@
    Experimental results on multiple datasets with different acceleration rates, sampling masks, and anatomical structures indicate that the proposed method can achieve better results without extra computational costs.
@@
    The proposed network design strategy can enhance the performance of other model-driven networks, making it a versatile approach for optimizing both single- and multi-prior based networks.
@@ 2 Problem Formulation
in sparsity-based algorithms. $\Phi$ is an appropriate transformation field for ensuring the sparsity of the raw data. We adopted an ISTA-based backbone [20] to solve (7), which iterates between the following two steps:
@@
For a low-rankness-based method, (6) can be expressed as
@@
$\hat{x}^{k}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x^{k-1}-y\|_{2}^{2}+\lambda_{2}\|\tau(x^{k-1})Q^{k-1}\|_{F}^{2}$ (13)
@@
$\hat{x}^{k}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x^{k-1}-y\|_{2}^{2}+\lambda_{2}\|\mathcal{D}(Q^{k-1})x^{k-1}\|_{F}^{2}$ (16)
Here we solve (16) by iterating the following two steps [55, 56, 44]:
@@ 3 Method
Figure 1: (a): the structure of the proposed network. (b): the structure of the sparsity based sub-network ($N_{s}$ in (a)). (c): the structure of the low-rank based sub-network ($N_{l}$ in (a)). (d): the structure of the attention modules (the yellow blocks in (a)), each of which include an information integration block (IIB), an attention block (AB), and a weighting block (WB). (e): the enriched information in IIBs. (f) the structure of the correction modules (the blue blocks in (a)), each of which includes a regularization block (RB) and a data consistency (DC) block.
@@ 3.1 Network Structure
The proposed network structure and its components are illustrated in Fig. 1. Fig. 1 (a) shows the overall network structure. In the network training process, the undersampled k-space data $y\in\mathbb{C}^{NC}$ as well as the corresponding sensitivity maps $\mathcal{S}\in\mathbb{C}^{NC}$ and masks $\mathcal{P}\in\mathbb{R}^{N}$ are fed to the network. MR images are stored using complex values, and some studies have proposed complex convolutions [13] to achieve better results; however, this is not our focus. Therefore, we leveraged the common method that two channels were used to store the real and imaginary parts in this study, and $y\in\mathbb{C}^{NC}$, $\mathcal{S}\in\mathbb{C}^{NC}$ were transferred to $y\in\mathbb{R}^{2NC}$, $\mathcal{S}\in\mathbb{R}^{2NC}$ before they are fed into the network. Besides, the coil-combined image $x^{0}$ is obtained from the zero-filled measurement $y$ and fed to the network. Then, several iteration blocks were stacked to form the main body, and each block is referred to as a cascade. Through stacked cascades, the network finally outputs the reconstructed MR images. Each cascade comprised three parts: MNs (green blocks), AMs (yellow blocks), and CMs (blue blocks). Fig. 1 (b)–(f) present the detailed structures of the modules, which are introduced in the following subsections.
3.2 Model-driven Subnetworks
@@
$N_{l}$ in each cascade is constructed by unrolling (17) and (18) into the network. For (17), essentially $(\mathcal{D}(Q))^{H}\mathcal{D}(Q)x$ can be regarded as a projection acting on $x$, which is determined by $Q$, and $Q$ at each iteration will also be updated according to the status of $x$. Therefore, we relaxed the constraint on (17) and used U-Net to learn the mapping directly, as shown in Fig. 1 (c). Although the constraint from the prior knowledge is relaxed, we can still exploit the complementary information from different subnetworks. Besides, the subnetwork gains enhanced flexibility to learn appropriate features from the data to deal with different input scenarios and improve its performance.
We used $N_{lu}$ to represent U-Net, and thus, we have
@@
3.3 Attention Modules
The AMs include the yellow blocks in Fig. 1; the details are shown in Fig. 1 (d). The first objective of an AM is to maximize the contributions of subnetworks in their respective areas of expertise. To achieve this, we design the RC maps to precisely weight the IRs produced by subnetworks in an element-wise manner. The RC essentially indicates the confidence level to accept the pixel values as the final predictions. Consequently, when the predicted value from one subnetwork matches the target closer than the other one, the RC value should be closer to one, signifying high confidence. We used $M^{k}_{r}$ to represent the RC map in the $k$-th cascade, and during the weighting process, we transferred $x_{s}^{k}$ and $x_{l}^{k}$ back to complex values (i.e., $x_{s}^{k},x_{l}^{k}\in\mathbb{R}^{2N}\rightarrow x_{s}^{k},x_{l}^{k}\in\mathbb{C}^{N}$) and then created a Hadamard product between $M^{k}_{r}$ and $x_{s}^{k}$. $x_{l}^{k}$ is similarly multiplied by $\textbf{J}-M^{k}_{r}$. We used $N_{a}$ to represent the weighting block (WB), and the weighted output is given by
$x^{k}_{a}=N_{a}(x_{s}^{k},x_{l}^{k})=x^{k}_{s}\odot M^{k}_{r}+x^{k}_{l}\odot(\textbf{J}-M^{k}_{r})$ (22)
where $x_{s}^{k}$ and $x_{l}^{k}$ are IRs produced by subnetworks in the $k$-th cascade, $x^{k}_{a}$ is the RC-weighted result and $\textbf{J}\in\mathbb{R}^{N}$ is a matrix filled with one.
However, there remains a flaw that hinders the proposed method from producing accurate results even with ideal (i.e., accurate) RC maps. In situations where the predicted values of a pixel from both subnetworks are simultaneously bigger or smaller than the target values, the RC-based weighting process fails to obtain more precise values, resulting in suboptimal convergence speed. In other words, even with ideal RC maps, the confidence level to accept $x^{k}_{a}$ as the final prediction is compromised in these pixels. To address the issue, we introduce OC maps, which learn the confidence of $x^{k}_{a}$, while CMs are designed to further optimize the low-OC areas. Consequently, the AMs are designed to simultaneously learn the RC and OC of the IRs in different cascades.
Empirically, learning from the richer features in subnetworks could improve the accuracy of RC and OC maps. However, this approach may introduce higher computational costs and potentially limit the versatility of the network. Therefore, we proposed the design of an information integration block (IIB) to learn the confidence maps from IRs across various subnetworks and cascades. By incorporating richer information from IR value maps across various stages, we can establish a more robust mapping to the target confidence maps, thereby enhancing the overall effectiveness of the network. The IIB in the $k$-th cascade is given as:
$x^{k}_{c}=\left\{\begin{array}[]{l}\mathcal{C}[x^{0},x_{s}^{k},x_{l}^{k}],\ \ \ \ \ \ \ k=1\\ \mathcal{C}[x^{0},x_{s}^{k-1},x_{l}^{k-1},x_{s}^{k},x_{l}^{k}],\ \ \ \ \ k\textgreater 1\\ \end{array}\right.$ (23)
where $x^{k}_{c}$ is the input of the attention blocks (ABs) and $\mathcal{C}[\cdot]$ denotes a concatenation operator on the channel dimension. Therefore, the input for the ABs consists of the coil-combined measurements $x^{0}$ and the IRs from the current and the preivous cascade (if they exists). This is aim to incorporate short and long-range information of the pixel values to build a stable mapping. It is worth emphasizing that the design of the input features is highly flexible, based on the fundamental concept that richer inputs generally yield more robust confidence maps. This will be further discussed in Section 4.3.
Then, $x_{c}^{k}$ is fed into an AB comprising two convolutional layers, each of which connects to an activation layer. Here we chose the ReLU and sigmoid functions, respectively. After the four-layer attention block, we obtained an two-channel attention map $M^{k}$, given as:
$M^{k}=\mathcal{C}[M^{k}_{r},M^{k}_{o}]=N_{AB}(x_{c}^{k})$ (24)
where $N_{AB}$ donotes the ABs. Each element in $M^{k}$ lies between zero and one. We separated $M^{k}$ into $M^{k}_{r}$ and $M^{k}_{o}$, where $M^{k}_{o}$ is the OC maps. The OC maps as well as the weighted IRs (i.e., $x^{k}_{a}$) are fed to the corrections modules.
3.4 Correction Modules
Through AMs, the pixels reconstructed by proper subnetworks can have larger weights, and the gap between intermediate outputs and targets can be narrowed. However, there may be instances where the intermediate values from both subnetworks are significantly bigger or smaller than the target values. In such cases, the RC-based weighting process encounters challenges in bringing the intermediate reconstructions closer to the target, leading to suboptimal convergence. Therefore, we proposed a correction module to further improve the network performance, especially on these pixels. Each correction module consists of a regularizer block (RB) and a data consistency (DC) block. The RB block is given as: